The temperature at which root mean square velocity of molecule of helium is equal to Root mean square velocity of hydrogen at N.T.P is

Grade:12th pass

1 Answers

Aashna
6 years ago
Answer will be 596K Root mean square velocity are given equal then Vrms=√3Rt/Mi.e Vrms ~T/MAs (Vrms)He =(Vrms)H2√T1/4=√T2/2As T2=298K (NTP)T1=596K